Two plastic containers have square bases, but the base of one container is 10 cm longer and wider than that of the other. Both containers are filled to a depth of 30 cm. Find the dimensions of the base of each container if the volume of water in the larger container exceeds that in the smaller container by 15,000 cm3 (cubic centimeters)
Answer by cleomenius(959)

(x+10)^2 * 30 = x^2 *30 + 15000 This is the equation to solve for x, the original length and width.
We can divide by 30 to simplify.
x^2 +20x+100 = x^2 + 500
20x + 100 = 500 subtract x^2 from each side.
20x = 400 subtract 100 from each side.
x = 20, the length and width, in cm
30cm = larger length and width.
20cm * 20cm * 30cm = 12000 cm cubed
30cm * 30cm * 30 cm = 27000 cm cubed
27000 cm cubed - 12000 cm cubed = 15000 cm cubed (check)
Cleomenius.